# Complementarity in Human-AI Collaboration: Concept, Sources, and Evidence

### v1

@@ 4 Experimental Studies
To demonstrate the proposed conceptualization’s value and application, and to further investigate human-AI decision-making in the presence of the identified sources of complementarity potential—information and capability asymmetry—we conduct two behavioral experiments. Specifically, we focus on a team setting in which a human decision-maker has access to AI advice and is subsequently responsible for making the final team decision based on his/her judgment and that of the AI (Green and Chen,, 2019). In this collaboration setup, the team decision either matches the individual human or AI decision—i.e., the human relies fully on his/her judgment or that of the AI—or it can be a function of the individual human and AI decision—i.e., an “integrated” decision that potentially differs from both individual ones. In the first experiment, we investigate the effect of information asymmetry on decision-making in the human-AI team in the form of additional contextual information only available to the human. In the second experiment, we investigate the effect of capability asymmetry on joint decision-making with different levels of diversity between the capabilities of humans and AI.
@@ 4.1 Experiment 1: The Effect of Information Asymmetry
In the first experiment, we apply the conceptualization developed in Section 3 to study the effect of information asymmetry between humans and AI as a relevant source of complementarity potential. More precisely, we create an intervention in which humans are given contextual information withheld from the AI to investigate whether and how this affects the final team decision and the realized complementarity effect.
@@ 4.1.1 Task and AI Model
We draw on a real estate appraisal task provided on the data science website kaggle.com (Kaggle,, 2019). Since housing is a basic need, and because it is ubiquitous in everyone’s life, all people to some degree have the ability to assess a house’s value on the basis of relevant factors such as size or appearance. The data set encompasses 15,474 houses and contains information about the street, city, number of bedrooms, number of bathrooms, and size (in square feet). In the data set, the house prices denote their listing price. The average house price is \$703,120, ranging between a minimum of \$195,000 and a maximum of \$2,000,000. An image of each house is also provided. In respect of the house price prediction task, we implement a random forest regression model as the AI model (Breiman,, 2001). We draw on the individual trees in the random forest to generate a predictive distribution for each instance and provide the 5% and 95% quantiles as indicators of the AI’s prediction uncertainty. We use 80% of the data as the training set and 20% as the test set. We train the random forest on the following features: the street, city, number of bedrooms, number of bathrooms, and square feet of the house. The house’s image is withheld from the AI model. In respect of the behavioral experiment, we focus on detached family houses in the test set, with the existing image providing a view of its exterior. From these, we randomly draw a hold-out set of 15 houses to serve as samples for our behavioral experiment. The AI model achieves a performance measured in terms of the mean absolute error (MAE) of \$163,080 regarding the hold-out set, which is representative of its performance on the entire test set. In respect of the condition with unique human contextual information (UHCI), we give humans an additional image of the house, which is likely to constitute valuable information. Humans are able to leverage their general understanding to form an overall assessment based on the house’s features, the visible surroundings, and its appearance. We conduct an initial pilot study to verify this assumption (Appendix B contains additional details).
@@ 4.1.2 Study Design
We conduct an online experiment with a between-subjects design. We recruit participants from prolific.com. The study includes two conditions and randomly assigns each participant to one of these conditions. We do not allow any repeated participation. Each participant passes the following steps (see Appendix B for additional details on the study design):
@@
Step 1: After accepting the task, participants are transferred to our experimental website. They are asked for their consent and to read the instructions. A control question initiates the study.
@@
Step 2: Since prior work highlights the importance of task training for the participants, we include a mandatory tutorial (Grootswagers,, 2020). In order to familiarize the participants with the task and data, both treatments receive an identical in-depth introduction to the data set, including summary statistics like the mean and the maximum and minimum house prices as reference points. Participants in the treatment without unique human contextual information (i.e., without UHCI) are only given the houses’ tabular data, while the participants in the treatment with unique human contextual information (i.e., with UHCI) are also given images of the house.
@@
Step 3: As part of the tutorial, we introduce the participants to the AI. We emphasize that the AI did not have access to the images during the training. We show the AI prediction in the context of the minimum and the maximum house prices. The participants also receive information about the AI’s uncertainty in the form of the 5% and 95% quantiles (see Figure 7). We explain the interpretation of the AI’s advice, including all the data points mentioned above. The participants are subsequently asked to answer a control question to verify their understanding.
@@
Step 4: The participants conduct two training task instances. For each instance, we initially let the participants to first provide a prediction, before we reveal the AI’s recommendation. They are then asked to adjust the AI’s prediction in the best possible way, constituting the joint human-AI team prediction. After each training example, the participants receive feedback in the form of the true house price. After completing the two training task instances, they are informed about the start of the study.
@@
Figure 7: An overview of the interfaces containing the information that the participants are given in the respective behavioral experiment’s treatments.
@@
Step 5: Each participant completes 15 house price prediction task instances, presented randomly in the same procedure as described in Step 4. During the task, the participants are not informed about the actual house price. Subsequently, we ask them to complete a questionnaire regarding qualitative feedback (Step 6) and demographic information (Step 7).
@@
The overall task lasts approximately 30 minutes. Before recruiting participants, we compute the required sample size in a power analysis using G*Power (Faul et al.,, 2007). Based on the pilot data, we expect a large effect ($d=$ 0.8). We refer to an alpha value of 0.05, while taking multiple testing into account in order to achieve a power of 0.8. This results in a total sample size of 86. Anticipating that some participants will fail the attention checks, we recruit a total of 120 participants (60 per condition). They receive a base payment of £5 and are additionally incentivized following the approach of Kvaløy et al., (2015), who show the benefits of combining non-monetary motivators, such as recognition, attention, and verbal feedback, with performance-based pay. We achieve this by adding motivational statements and by giving the top 10% participants an additional pound. Note that the two training task instances are not included in the final evaluation. To ensure the quality of the collected data, we remove those participants whose entered prices exceed the communicated maximum house price of \$2,000,000 in the data set. We also identify outliers for removal by using the median absolute deviation (Leys et al.,, 2013; Rousseeuw and Croux,, 1993). After applying these criteria, we continue with the data of 101 participants across both conditions—53 in the treatment without UHCI and 48 in that with UHCI (see Appendix B for additional details about the participants).
@@ 4.1.3 Evaluation Measures
For each participant, we measure the loss of the human ($l_{H}$), the AI ($l_{AI}$), and the team decision ($l_{I}$) as the absolute error, and calculate the average over all task instances to receive the human ($L_{H}$), the AI ($L_{AI}$), and the team performance ($L_{I}$) corresponding to the mean absolute error (MAE). Furthermore, we calculate the complementarity potential’s and complementarity effect’s respective components as defined in Section 3. Finally, for each measure, we calculate the average over all the participants.
@@ 4.1.4 Results
In this section, we analyze the impact of unique human contextual information on the team performance, the complementarity potential, and the complementarity effect. We evaluate the results’ significance by using the Student’s T-test and the Mann-Whitney U-test, depending on whether the prerequisites have been fulfilled. We apply the Bonferroni correction and adjust the p-values accordingly. First, we focus on the impact of contextual information on performance, followed by an in-depth analysis of its effect on the complementarity potential and on its constituting components.
@@
First, we analyze the inherent complementarity potential (${CP}^{inh}$). We observe a significant increase due to the unique human contextual information. In the condition without UHCI, the ${CP}^{inh}$ is \$42,995, and increases to \$61,970 in the condition with UHCI ($d=$ 1.05, $p<$ 0.001, two-tailed Mann-Whitney U test). This finding can be interpreted that the images indeed contain useful contextual information for humans, which the AI cannot access.
Next, we calculate the collaborative complementarity potential (${CP}^{coll}$). Whereas in the condition without UHCI, the ${CP}^{coll}$ results in \$120,085, in the condition with UHCI it amounts to \$101,110. The difference is statistically significant ($d=$ 1.05, $p<$ 0.001, two-tailed Mann-Whitney U test). Finally, since the participants in both conditions work with the same AI, which has an overall better individual performance, the $CP$ (i.e., the sum of the inherent and collaborative component) is \$163,080 in both conditions. As ${CP}^{inh}$ increases due to UHCI, ${CP}^{coll}$ decreases because the $CP$ remains constant.
@@
Next, we focus on the realized complementarity potential, i.e., the complementarity effect ($CE$). We find a significant difference between the inherent complementarity effect (${CE}^{inh}$) in both conditions (without UHCI: \$14,468; with UHCI: \$27,860; $d=$ 0.87, $p<$ 0.001, two-tailed Mann-Whitney U test), which highlights contextual information’s potential. This absolute increase might be due to an increase in the complementarity potential and/or an improvement in the integration of both team members’ predictions through the human. In this context, in order to investigate this further, we also calculate the inherent complementarity effect’s ($\frac{{CE}^{inh}}{{CP}^{inh}}$) relative amount. This analysis reveals that unique human contextual information not only enhances the theoretically available inherent complementarity potential, but that the participants could also use significantly more of it (without UHCI: 34%; with UHCI: 45%; $d=$ 0.82, $p<$ 0.001, two-tailed Mann-Whitney U test). This is an interesting result, as having more information available than the AI might also have detrimental psychological effects. Jussupow et al., (2020), for example, find that perceived AI capabilities and expertise influence aversion towards AI. This could lead to humans taking less account of the AI’s suggestions when making their final decision (Longoni et al.,, 2019; Mahmud et al.,, 2022; Sieck and Arkes,, 2005), which could result in $CTP$ not being achieved. However, our results show that unique human contextual information does not only increase the potential, but also the integration’s overall effectiveness.
Figure 9: Summary of the real estate appraisal experiment’s results.
Next, we analyze unique human contextual information’s impact on the collaborative complementarity effect (${CE}^{coll}$). We do not find a significant difference between the two treatments (${CE}^{coll}$: without UHCI: \$-11,483; with UHCI: \$-12,789; $d=$ 0.08, $p=$ 1.0, two-tailed Mann-Whitney U test). In our experiment, it seems intuitive for the ${CE}^{coll}$ to yield a negative value. Given that the AI in our setup outperforms its human team member, a positive ${CE}^{coll}$ could only occur on task instances where the team loss is even lower than that of the AI and the human (see Figure 4). Conversely, each instance where the human underperforms in comparison to the AI and fails to adopt the AI’s decision contributes negatively to the ${CE}^{coll}$. Since humans tend to choose decisions between two boundaries (in our case their own and the AI’s decision), our setup naturally fosters a negatively realized collaborative complementarity potential. See Appendix B for further results, including a performance analysis of each house.
@@ 4.2 Experiment 2: The Effect of Capability Asymmetry
In the second behavioral experiment, we apply the conceptualization to study the effect of capability asymmetry between humans and AI as another relevant source of complementarity potential. In detail, starting with a “baseline” AI, we create an intervention in which we increase the asymmetry between the AI’s and the humans’ capabilities while keeping its overall performance constant. This means the second AI tends to make correct decision suggestions for task instances that tend to be more difficult for humans (“complementary” AI).
@@ 4.2.1 Task and AI Model
To investigate the effect of capability asymmetry on the human-AI team performance in decision-making, we choose the image recognition context. Research has demonstrated that humans and AI tend to make different errors on image classification tasks (Fügener et al., 2021a, ; Steyvers et al.,, 2022). Specifically, an AI based on deep convolutional neural networks tends to infer classification decisions differently than humans do (Geirhos et al.,, 2020, 2021). We could therefore expect the AI model to classify certain images more accurately than humans and vice versa, thereby creating inherent complementarity potential. However, it remains unclear whether this naturally existing potential could be sufficiently realized when humans incorporate the AI decision into a final team decision, and whether its increase in the intervention affects the realization.
@@
In order to undertake the experiment, we draw on the image data set that Steyvers et al., (2022) provided. The data set comprises 1,200 images distributed evenly across 16 classes (e.g., airplane, dog, or car). It was curated on the basis of the ImageNet Large Scale Visual Recognition Challenge (ILSVRC) 2012 database (Russakovsky et al.,, 2015). To increase the task difficulty for humans and the AI, the authors applied phase noise distortion at each spatial frequency, which was uniformly distributed in the interval $[-\omega,\omega]$ with $\omega=$ 110. Despite the heightened difficulty level, both humans and AI can attain comparable performance on the task. In addition to ground truth labels, the data set also contains multiple human predictions for each image provided by crowd workers, allowing us to infer a proxy for human classification difficulty. Images with a high disagreement in respect of multiple human predictions indicate a higher level of difficulty for humans.
@@
We implement the AI model as a convolutional neural network, more precisely, as a DenseNet161 (Huang et al.,, 2017), pre-trained on ImageNet. We partition the data set into a training (60%), validation (20%), and test set (20%). In the baseline condition, we fine-tune the AI model on the distorted images over 100 epochs, applying early stopping on the validation loss. We use SGD as an optimizer with a learning rate of $1\cdot{10}^{-4}$, a weight decay of $5\cdot{10}^{-4}$, a cosine annealing learning rate scheduler, and a batch size of 32. The AI model achieves a classification error of 26.66% on the test set.
@@
In the intervention, we create an alternative AI model that makes erroneous decisions for different instances. We fine-tune the DenseNet161 model exactly as in the baseline condition, but, for each image in the training set, also incorporate a human prediction as an additional label in the training process in order to incentivize the AI model to learn to correctly classify the images that tend to be more difficult for humans (Hemmer et al.,, 2022; Madras et al.,, 2018; Wilder et al.,, 2020). See Appendix C for additional implementation details of this approach. Even though the AI model has a slightly higher classification error of 33.75%, this approach results in more asymmetric, i.e. non-overlapping, capabilities between the humans and the AI model. We select 15 images from the test set for the experiment, such that both AI models exhibit the same performance on the test set (26.66%), while considering non-overlapping errors between both AI models.

### v2

@@ 4 Experimental Studies
To demonstrate the proposed conceptualization’s value and application, and to further investigate human-AI decision-making in the presence of the identified sources of complementarity—information and capability asymmetry—we conducted two behavioral experiments. Specifically, we focused on a team setting in which a human decision-maker has access to AI advice and is subsequently responsible for making the final team decision based on his/her judgment and that of the AI (Green and Chen, 2019). In this collaboration setup, the team decision either matches the individual human or AI decision—i.e., the human relies fully on his/her judgment or that of the AI—or it can be a function of the individual human and AI decision—i.e., an “integrated” decision that potentially differs from both individual ones. In the first experiment, we investigated the effect of information asymmetry on decision-making in the human-AI team in the form of additional contextual information only available to the human. In the second experiment, we investigated the effect of capability asymmetry on joint decision-making with different levels of diversity between the capabilities of humans and AI.
@@ 4.1 Experiment 1: The Effect of Information Asymmetry
In the first experiment, we applied the conceptualization developed in Section 3 to study the effect of information asymmetry between humans and AI as a relevant source of complementarity. More precisely, we created an intervention in which humans are given contextual information withheld from the AI to investigate whether and how this affects the final team decision and the realized complementarity effect.
@@ 4.1.1 Task and AI Model
We drew on a real estate appraisal task provided on the data science website kaggle.com (Kaggle, 2019). Since housing is a basic need, and because it is ubiquitous in everyone’s life, all people to some degree have the ability to assess a house’s value on the basis of relevant factors such as size or appearance. The data set encompasses 15,474 houses and contains information about the street, city, number of bedrooms, number of bathrooms, and size (in square feet). In the data set, the house prices denote their listing price. The average house price is \$703,120, ranging between a minimum of \$195,000 and a maximum of \$2,000,000. An image of each house is also provided.
In respect of the house price prediction task, we implemented a random forest regression model as the AI model (Breiman, 2001). We drew on the individual trees in the random forest to generate a predictive distribution for each instance and provided the 5% and 95% quantiles as indicators of the AI model’s prediction uncertainty. We used 80% of the data as the training set and 20% as the test set. We trained the random forest on the following features: the street, city, number of bedrooms, number of bathrooms, and square feet of the house. The house’s image was withheld from the AI model.
In respect of the behavioral experiment, we focused on detached family houses in the test set, with the existing image providing a view of its exterior. From these, we randomly drew a hold-out set of 15 houses to serve as samples for our behavioral experiment. The AI model achieves a performance measured in terms of the mean absolute error (MAE) of \$163,080 regarding the hold-out set, which is representative of its performance on the entire test set. In respect of the condition with unique human contextual information (UHCI), we gave humans an additional image of the house, which is likely to constitute valuable information. Humans are able to leverage their general understanding to form an overall assessment based on the house’s features, the visible surroundings, and its appearance. We conducted an initial pilot study to verify this assumption (Appendix B.1 contains additional details).
@@ 4.1.2 Study Design
We conducted an online experiment with a between-subject design. We recruited participants from prolific.com. The study included two conditions and randomly assigned each participant to one of these conditions. We did not allow any repeated participation. Each participant passed the following steps (see Figure B.1 in Appendix B.2 for additional details):
@@
Step 1: After accepting the task, participants were transferred to our experimental website. They were asked for their consent and to read the instructions. A control question initiated the study.
@@
Step 2: Since prior work highlights the importance of task training for the participants, we included a mandatory tutorial (Grootswagers, 2020). In order to familiarize the participants with the task and data, both treatments received an identical in-depth introduction to the data set, including summary statistics like the mean and the maximum and minimum house prices as reference points (see Figure B.2 in Appendix B.2). Participants in the treatment without unique human contextual information (i.e., without UHCI) were only given the houses’ tabular data, while the participants in the treatment with unique human contextual information (i.e., with UHCI) were also given images of the house.
@@
Step 3: As part of the tutorial, we introduced the participants to the AI. We emphasized that the AI did not have access to the images during the training. We showed the AI prediction in the context of the minimum and the maximum house prices. The participants also received information about the AI’s uncertainty in the form of the 5% and 95% quantiles (see Figure 7). We explained the interpretation of the AI’s advice, including all the data points mentioned above (see Figure B.3 in Appendix B.2). The participants were subsequently asked to answer a control question to verify their understanding.
@@
Step 4: The participants conducted two training task instances. For each instance, we initially asked the participants to first provide a prediction, before we revealed the AI’s recommendation. They were then asked to adjust the AI’s prediction in the best possible way, constituting the joint human-AI team prediction. After each training example, the participants received feedback in the form of the true house price (see Figure B.4 in Appendix B.2 for an exemplary training task instance). After completing the two training task instances, they were informed about the start of the study.
@@
Figure 7: An overview of the interfaces containing the information that the participants were given in the respective behavioral experiment’s treatments.
@@
Step 5: Each participant completed 15 house price task instances, presented randomly in the same procedure as described in Step 4 (see Figure B.5 in Appendix B.2 for an exemplary task instance). During the task, the participants were not informed about the actual house price. Subsequently, we asked them to complete a questionnaire regarding qualitative feedback (Step 6) and demographic information (Step 7).
@@
The overall task lasted approximately 30 minutes. Before recruiting participants, we computed the required sample size in a power analysis using G*Power (Faul et al., 2007). Based on the pilot data, we expected a large effect ($d=$ 0.8). We referred to an alpha value of 0.05, while taking multiple testing into account in order to achieve a power of 0.8. This resulted in a total sample size of 86. Anticipating that some participants will fail the attention checks, we recruited a total of 120 participants (60 per condition). They received a base payment of £5 and are additionally incentivized following the approach of Kvaløy et al. (2015), who show the benefits of combining non-monetary motivators, such as recognition, attention, and verbal feedback, with performance-based pay. We achieved this by adding motivational statements and by giving the top 10% participants an additional pound. Note that the two training tasks are not included in the final evaluation. To ensure the quality of the collected data, we removed those participants whose entered prices exceed the communicated maximum house price of \$2,000,000 in the data set. We also identified outliers for removal by using the median absolute deviation (Leys et al., 2013; Rousseeuw and Croux, 1993). After applying these criteria, we continued with the data of 101 participants across both conditions—53 in the treatment without UHCI and 48 in that with UHCI (see Table B.1 in Appendix B.3 for additional details about the participants).
@@ 4.1.3 Evaluation Measures
For each participant, we measured the loss of the human ($l_{H}$), the AI ($l_{AI}$), and the team decision ($l_{I}$) as the absolute error, and calculated the average over all task instances to receive the human ($L_{H}$), the AI ($L_{AI}$), and the team performance ($L_{I}$) corresponding to the mean absolute error (MAE). Furthermore, we calculated the complementarity potential’s and complementarity effect’s respective components as defined in Section 3. Finally, for each measure, we calculated the average over all the participants.
@@ 4.1.4 Results
In this section, we analyze the impact of unique human contextual information on the team performance, the complementarity potential, and the complementarity effect. We evaluate the results’ significance by using the Student’s T-test and the Mann-Whitney U-test, depending on whether the prerequisites have been fulfilled. We apply the Bonferroni correction and adjust the p-values accordingly. First, we focus on the impact of contextual information on performance, followed by an in-depth analysis of its impact on complementarity potential and effect.
@@
First, we analyze the inherent complementarity potential (${CP}^{inh}$). We observe a significant increase due to the unique human contextual information. In the condition without UHCI, the ${CP}^{inh}$ is \$42,995, and increases to \$61,970 in the condition with UHCI ($d=$ 1.05, $p<$ 0.001, two-tailed Mann-Whitney U test). This finding can be interpreted that the images indeed contain useful contextual information for humans, which the AI cannot access, resulting in fewer shared errors that humans and AI make individually.
Next, we calculate the collaborative complementarity potential (${CP}^{coll}$). Whereas in the condition without UHCI, the ${CP}^{coll}$ results in \$120,085, in the condition with UHCI it amounts to \$101,110. The difference is statistically significant ($d=$ 1.05, $p<$ 0.001, two-tailed Mann-Whitney U test). Since the participants in both conditions work with the same AI model, which has an overall better individual performance, the $CP$ (i.e., the sum of the inherent and collaborative component) is \$163,080 in both conditions. As ${CP}^{inh}$ increases due to UHCI, ${CP}^{coll}$ decreases because the $CP$ remains constant.
@@
Then, we focus on the realized complementarity potential, i.e., the complementarity effect ($CE$). We find a significant difference between the inherent complementarity effect (${CE}^{inh}$) in both conditions (without UHCI: \$14,468; with UHCI: \$27,860; $d=$ 0.87, $p<$ 0.001, two-tailed Mann-Whitney U test), which highlights contextual information’s potential. This absolute increase might be due to an increase in inherent complementarity potential and/or an improvement in the integration of both team members’ predictions through the human. In order to investigate this further, we also calculate the inherent complementarity effect’s ($\frac{{CE}^{inh}}{{CP}^{inh}}$) relative amount. This analysis reveals that unique human contextual information not only enhances the theoretically available inherent complementarity potential, but that the participants could also use significantly more of it (without UHCI: 34%; with UHCI: 45%; $d=$ 0.82, $p<$ 0.001, two-tailed Mann-Whitney U test). This is an interesting result, as having more information available than the AI might also have detrimental psychological effects. Jussupow et al. (2020), for example, find that perceived AI abilities and expertise influence aversion towards AI. This could lead to humans taking less account of the AI’s suggestions when making their final decision (Longoni et al., 2019; Mahmud et al., 2022; Sieck and Arkes, 2005), which could result in CTP not being achieved. However, our results show that unique human contextual information does not only increase the potential, but also the integration’s overall effectiveness.
Figure 9: Result summary of the real estate appraisal experiment.
Next, we analyze unique human contextual information’s impact on the collaborative complementarity effect (${CE}^{coll}$). We do not find a significant difference between the two treatments (${CE}^{coll}$: without UHCI: \$-11,483; with UHCI: \$-12,789; $d=$ 0.08, $p=$ 1.0, two-tailed Mann-Whitney U test). In our experiment, it seems intuitive for the ${CE}^{coll}$ to yield a negative value. Given that the AI in our setup outperforms its human team member, a positive ${CE}^{coll}$ could only occur on task instances where the team loss is even lower than that of the AI and the human (see Figure 5). Conversely, each instance where the human underperforms in comparison to the AI and fails to improve the AI’s decision contributes negatively to the ${CE}^{coll}$. Since humans tend to choose decisions between two boundaries (in our case their own and the AI’s decision), our setup naturally fosters a negative collaborative complementarity effect. Nevertheless, investigating the individual performance of humans and AI as well as the team performance for each house separately reveals that the human-AI team can derive team decisions for task instances 1, 4, and 13 that are, on average over all participants, more accurate compared to the respective individual decisions of human and AI (see Table B.2 in Appendix B.4). This demonstrates the occurrence of a positive collaborative complementarity effect for these three task instances. See Appendix B.4 for further results regarding a performance analysis of each house.
@@ 4.2 Experiment 2: The Effect of Capability Asymmetry
In the second behavioral experiment, we applied the conceptualization to study the effect of capability asymmetry between humans and AI as another relevant source of complementarity. In detail, starting with a “baseline” AI, we created an intervention in which we increase the asymmetry between the AI’s and the humans’ capabilities while keeping its overall performance constant. This means the second AI tends to make correct decision suggestions for task instances that tend to be more difficult for humans (“complementary” AI).
@@ 4.2.1 Task and AI Model
To investigate the effect of capability asymmetry on the human-AI team performance in decision-making, we chose the image recognition context. Research has demonstrated that humans and AI tend to make different errors on image classification tasks (Fügener et al., 2021; Steyvers et al., 2022). Specifically, an AI model based on deep convolutional neural networks tends to infer classification decisions differently than humans do (Geirhos et al., 2020, 2021). We could therefore expect the AI model to classify certain images more accurately than humans and vice versa, thereby creating inherent complementarity potential. However, it remains unclear whether this naturally existing potential could be sufficiently realized when humans incorporate the AI decision into a final team decision, and whether its increase in the intervention affects the realization.
@@
In order to undertake the experiment, we drew on the image data set that Steyvers et al. (2022) provided. The data set comprises 1,200 images distributed evenly across 16 classes (e.g., airplane, dog, or car). It was curated on the basis of the ImageNet Large Scale Visual Recognition Challenge (ILSVRC) 2012 database (Russakovsky et al., 2015). To increase the task difficulty for humans and the AI, the authors applied phase noise distortion at each spatial frequency, which was uniformly distributed in the interval $[-\omega,\omega]$ with $\omega=$ 110. Despite the heightened difficulty level, both humans and AI can attain comparable performance on the task. In addition to ground truth labels, the data set also contains multiple human predictions for each image provided by crowd workers, allowing us to infer a proxy for human classification difficulty. Images with a high disagreement in respect of multiple human predictions indicate a higher level of difficulty for humans.
@@
We implemented the AI model as a convolutional neural network, more precisely, as a DenseNet161 (Huang et al., 2017), pre-trained on ImageNet. We partitioned the data set into a training (60%), validation (20%), and test set (20%). In the baseline condition, we fine-tuned the AI model on the distorted images over 100 epochs, applying early stopping on the validation loss. We used SGD as an optimizer with a learning rate of $1\cdot{10}^{-4}$, a weight decay of $5\cdot{10}^{-4}$, a cosine annealing learning rate scheduler, and a batch size of 32. The AI model achieves a classification error of 26.66% on the test set.
@@
In the intervention, we created an alternative AI model that makes erroneous decisions for different instances. We fine-tuned the DenseNet161 model exactly as in the baseline condition, but, for each image in the training set, also incorporated a human prediction as an additional label in the training process in order to incentivize the AI model to learn to correctly classify the images that tend to be more difficult for humans (Hemmer et al., 2022; Madras et al., 2018; Wilder et al., 2020). See Appendix C.1 for additional implementation details of this approach. Even though the AI model has a slightly higher classification error of 33.75%, this approach results in higher capability asymmetry, i.e., non-overlapping capabilities between the humans and the AI model. We selected 15 images from the test set for the experiment, such that both AI models exhibit the same performance on the test set (26.66%), while considering non-overlapping errors between both AI models.